OpenStudy (anonymous):

determine the best method to solve each system of equation.then solve the system. 1.x+2y=3; x+y=1 2.m+6n=-8; m=2n+8 3.a-b=6; a=2b+7 4.4x+y=15; -x-3x=-12 5.3c-d=14; c-d=2 6.x+2y=-9; y=4x

OpenStudy (anonymous):

x+2y=3 x+ y=1 Subtract: -> y = 2 => x = 1 -2 = -1 Thus ( -1, 1)

OpenStudy (anonymous):

m+6n = -8 m -2n = 8 Subtract: -> 8n = -16 => n = -2, => m = 8 + 2 ( -2) = 4 Thus ( 4, -2)

OpenStudy (anonymous):

a-b =6; a-2b= 7 Subtract: -> b = -1 => a = 6 -1 = 5 Thus (5, -1)

OpenStudy (anonymous):

4x+y= 15; x = -3y + 12 Substitute: -> 4 (-3y + 12) +y = 15 -12y +48 + y = 15 => y = ( 15 - 48)/ -11 = 3 then x = -3 * 3 + 12 = 3 Thus ( 3,3)

OpenStudy (anonymous):

5. elimination method 6. substitution You should get the concept by now. It's good practice to arrange these equations vertically.